MODE SPECIFIC DYNAMICS IN THE PREDISSOCIATED, QUASILINEAR B^{1}A′ STATE OF CHF PROBED BY OPTICAL-OPTICAL DOUBLE RESONANCE SPECTROSCOPY

\underline{C. TAO}, C. MUKARAKATE AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233; T. W. SCHMIDT AND S. H. KABLE, School of Chemistry, University of Sydney, NSW 2006, Australia.

We have recently observed transitions to the predissociated, quasilinear B^{1}A′ state of a halocarbene, CHF, using a fluorescence dip detected optical-optical double resonance technique via the A^{1}A′′ state.^{a} By exciting selected rotational levels in intermediate states belonging to the progressions 2_{0}^{n}, 1_{0}^{1}2_{0}^{n}, and 2_{0}^{n}3_{0}^{1}, a variety of B^{1}A′ state levels have been observed, extending to an energy of 7000 cm^{−1} above the B^{1}A′ state origin. In this talk, we will focus on the dynamics of the B^{1}A′ state. All of the observed lines are predissociated, as evidenced by Lorentzian lineshapes, and the linewidths increase with increasing energy. A pronounced mode specificity is observed; levels containing CF stretching excitation dissociate more rapidly than nearly isoenergetic bending levels. The implications of these results for the dissociation mechanism will be emphasized.

aC. Tao, S. A. Reid, T. W. Schmidt, and S. H. Kable, J. Chem. Phys. 125, 051105 (2007).